OpenStudy (anonymous):

How can you make 0.05 into a fraction? please help!

OpenStudy (anonymous):

5/100

OpenStudy (anonymous):

5 out of 100 .5 = 1/2 .05 =1/20 .05 x 20 = 1

OpenStudy (anonymous):

thnx! :)